OpenStudy (konradzuse):

Find the equation for the tangent line of the graph y = sinh(1-x^2) for points (1,0)

OpenStudy (amistre64):

ewww, sinh

OpenStudy (amistre64):

we got a formula for that one?

OpenStudy (amistre64):

sinh(x) = (e^-x - e^ix)/2 or something like thjat

OpenStudy (amistre64):

lol, that was simple enough; D[sinh] = cosh

OpenStudy (konradzuse):

haha yeah sinh = cosh sinh = e^x - e^-x/2

OpenStudy (amistre64):

the eq of the tan line is just: \[tan_y=f'(1)x-f'(1)1+0\]

OpenStudy (amistre64):

\[f' = -2x\ cosh(1-x^2)\]

OpenStudy (amistre64):

f'(1) = -2 cosh(0) whatever that maybe

OpenStudy (konradzuse):

so all we do is plug in the points 1 and 0... after taking the derivative?

OpenStudy (amistre64):

well, the point is what we use to calibrate the line itself; and since x=1 is our value for x, we plug that into the f'(x) to get the value of the slope at (1,0)

OpenStudy (amistre64):

so yeah, slope = f' slope at x=1 is f'(1)

OpenStudy (konradzuse):

okay... So it could be any number, it just needs to create the line it is asked?

OpenStudy (amistre64):

correct

OpenStudy (amistre64):

If they had given a different point, say (3,-7) then slope = f'(3) and the rest of the line equation from algrebra forms as it forms

OpenStudy (konradzuse):

okay, so we only use the x value then, ic. So why is the formula f(1)x - f(1)1 - 0? Where do we use that?

OpenStudy (amistre64):

recall the point slope form of a line? y -Py = m(x-Px) y = mx -mPx + Py ; (Px,Py)

OpenStudy (amistre64):

m=f', (Px, Py) is given as the point

OpenStudy (konradzuse):

ah okay so we need to find m, and then have our equation. so it's y - 0= m (x-1)?

OpenStudy (amistre64):

correct; and m = f'(1) in this case

OpenStudy (amistre64):

derivative is equation for slope at any given point

OpenStudy (konradzuse):

okay so for the 3,7 point f(3) = m?

OpenStudy (konradzuse):

f'(3)

OpenStudy (amistre64):

correct

OpenStudy (konradzuse):

okay so in this case it would be y-0 = cosh(0) [1] (x-1) y = 1(x-1) = 1x-1

OpenStudy (amistre64):

f' = -2x cosh(1-x^2) ; when x=1; f'(1) = -2cosh(0)

OpenStudy (amistre64):

cosh(0) = 1 so f'(1) = -2

OpenStudy (konradzuse):

oh boo I thought I was cool and didn't check the answer and guessed haha... :(

OpenStudy (konradzuse):

okay so y = -2x +2

OpenStudy (amistre64):

tany = -2x+2(1)+0 tany = -2x+2 ; correct

OpenStudy (konradzuse):

ah so that's where that formula comes in :)

OpenStudy (amistre64):

yeah, i find it simper to work than the usual point slope form that you have to then work around to get it into this form anyways

OpenStudy (konradzuse):

as long as you can test it :)

OpenStudy (konradzuse):

tany=f′(1)x−f′(1)1+0 One o more thing on this, if it was the point (3,7) would it be tany=f′(3)x−f′(3)1+0

OpenStudy (amistre64):

no; IF the point had been (3,7) we would have used this point instead. tany = f'(3)x -f'(3)3 + 7

OpenStudy (konradzuse):

so f'(x)x - f'(x)x + 7?

OpenStudy (konradzuse):

except the 2nd x is a constant.

OpenStudy (konradzuse):

+ y

OpenStudy (konradzuse):

GRR this test isn't going to go well :'(

OpenStudy (amistre64):

close, except for the equation of the tangent line use a specific value associated with the derivative; so instead of confusing xs, I use Px and Py for the pointX component and pointY component tany = f'(Px)x - f'(Px)Px + Py

OpenStudy (amistre64):

you might be used to subscripts for your given point and generic for the unknowns \[y_{tan}=f'(x_o)x-f'(x_o)+y_0\]

OpenStudy (amistre64):

or even: \[y_{tan}=f'(c)x-f'(c)c+f(c)\]

OpenStudy (amistre64):

such that the point given is: (c,f(c))

OpenStudy (konradzuse):

tany = f'(Px)x - f'(Px)Px + Py works :)

OpenStudy (amistre64):